OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 8x - y = -9 4x - 3y = -22

OpenStudy (***[isuru]***):

now consider the first equation 8x - y = -9 make y the subject y = 8x +9 now plug this value in the second equation u got 4x - 3y = -22 4x - 3( 8x + 9) = -22 now simplify this.. it will give u the new equation the question asks.. can u do that :) ?

OpenStudy (lukebluefive):

Following the instructions from the problem, take the first equation and get "y" on a side by itself: \[8x - y = -9\] Add y to both sides: \[8x = y - 9\] Add 9 to both sides: \[8x + 9 = y\] Rotate the equation: \[y = 8x + 9\] Then do what Isuru says.